# Generative Refocusing: Flexible Defocus Control from a Single Image

### v2

@@ 3.3 Bokeh-Shape Aware Synthesis
Table 2: Defocus deblurring benchmark on RealDOF [36] and DPDD [1] datasets. We evaluate the task of recovering all-in-focus images from defocused-blur inputs. Our method, GenRefocus, uses the DeblurNet module for this task. We report both reference-based metrics (LPIPS, FID) and no-reference perceptual quality metrics (CLIP-IQA, MANIQA, MUSIQ). Best, second best, and third best results are highlighted. GenRefocus outperforms all competing approaches across both benchmarks and all metrics, demonstrating the effectiveness of our diffusion-based deblurring with pre-deblur conditioning.
@@
Method LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
Input 0.5241 92.93 0.3562 0.2213 28.7087 0.3485 88.85 0.4337 0.3325 45.5376
GT 0.4233 0.2675 42.5777 0.4563 0.3294 47.8102
AIFNet [55] 0.4132 72.38 0.3814 0.2536 35.7121 0.3090 79.84 0.4359 0.3371 46.9558
IFANet [36] 0.2994 33.21 0.4088 0.2609 39.5949 0.2301 42.60 0.4472 0.3276 47.2784
DRBNet [56] 0.2550 29.89 0.3889 0.2609 38.9014 0.1819 39.37 0.4142 0.3254 46.7229
Restormer [84] 0.2863 29.80 0.4062 0.2642 40.8355 0.1762 33.45 0.4332 0.3307 47.9872
INIKNet [51] 0.2851 32.18 0.3984 0.2580 38.9959 0.1860 36.63 0.4290 0.3265 47.1961
GenRefocus (Ours) 0.2356 24.73 0.4575 0.2886 43.7042 0.1598 33.08 0.4619 0.3411 49.3019
Figure 4: Qualitative comparison on defocus deblurring. Visual results on (a) RealDOF [36] and (b) DPDD [1] datasets. Blue boxes on the left indicate cropped regions shown in detail. Our DeblurNet faithfully recovers fine text details (“NEW YORK”) in (a) where competing methods produce blurry or distorted results. In the challenging example (b) with severe defocus blur, other methods either fail to recover structure (AIFNet [55], DRBNet [56], INIKNet [51]) or introduce artifacts in the background (IFANet [36], Restormer [84]), while our method reconstructs geometrically consistent, visually compelling content by leveraging diffusion priors guided by the pre-deblurred input.
Simulator and Data.
Real photographs exhibiting diverse bokeh shapes are rare, and paired AIF–bokeh examples are even scarcer. We thus rely on simulation. A key observation is that when an all-in-focus (AIF) image lacks point-light stimuli, the simulated bokeh carries weak shape cues, making the model reluctant to learn shape-conditioned responses. To address this, we synthesize a PointLight-1K AIF set designed to reveal aperture shape. We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $S_{1}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator renders
$I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,S_{1},K,s\big),$ (3)
@@
Shape-conditioned fine-tuning.
Starting from our base model trained without shape control, we freeze all original LoRA weights and introduce another conditioning branch that ingests a shape exemplar $s$. Only this branch is fine-tuned while the backbone remains fixed, ensuring that shape edits do not regress the learned bokeh synthesis.
@@ 4.1 Setup
Backbone and training details.
Our base generative backbone is FLUX-1-dev [61], fine-tuned via LoRA [28] with a multi-condition scheme following [66]. DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64.
@@
Both models are trained with per-GPU batch size $1$ and gradient accumulation over $8$ steps on $4{\times}$RTX A6000 GPUs. DeblurNet is trained for $60\mathrm{K}$ steps. BokehNet is trained in two stages: synthetic pair training for $40\mathrm{K}$ steps and Real unpair for another $40\mathrm{K}$ steps.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT BokehMe [48] Bokehlicious [60] BokehDiff [89] Ours
Figure 5: Qualitative comparison on bokeh synthesis benchmark. Results on LF-Bokeh with zoomed patches (blue and orange boxes) highlighting detail quality. Our BokehNet synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling. Baselines show various artifacts: BokehMe [48] exhibits simulator bias, Bokehlicious [60] over-smooths details, and BokehDiff [89] produces inconsistent defocus. Our semi-supervised training on real bokeh images enables the capture of authentic lens characteristics.
Datasets.
We summarize the training data for the two modules below.
DeblurNet. We train on DPDD [1] together with a curated subset of RealBokeh_3MP [60], totaling approximately 4.5K image pairs.
BokehNet (synthetic pairs). We use AIF images from [82] and a subset of EBB! [30], and procedurally render defocus under varying aperture sizes/shapes and focus points to generate about 70K synthetic pairs.
BokehNet (real supervision). We incorporate (i) an unpaired real-bokeh set collected from Flickr by Fortes et al. [19], and (ii) real AIF–bokeh pairs with incomplete metadata from RealBokeh_3MP [60] and LFDOF [55], contributing approximately 27K real examples in total. All datasets are used strictly for research and are split so that no scene overlaps across training and evaluation.
Table 3: Bokeh synthesis benchmark. Evaluation on LF-Bokeh (200 light-field images with multiple focus planes). Following [89], we use per-image binary search for optimal $K$. Bokehlicious supports only aperture size control. Our semi-supervised training on real bokeh data outperforms all baselines.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
BokehMe [48] 0.1228 0.0744 0.9511
Bokehlicious [60] 0.1799 0.1062 0.9304
BokehDiff [89] 0.1708 0.0933 0.9192
GenRefocus (Ours) 0.1047 0.0611 0.9570
Benchmarks.
We evaluate three tasks: (i) defocus deblurring on DPDD and RealDOF; (ii) bokeh synthesis on our new LF-Bokeh benchmark, comprising $200$ images with multiple focus planes synthesized from [53] and [15] light-field captures; and (iii) refocusing on LF-Refocus, built by pairing same-scene images within LF-Bokeh to form $400$ source–target pairs.
Inference-time resolution and aspect ratio.
We adopt a tiling strategy inspired by  [5] at evaluation time: we split the input image into overlapping patches and perform tile-wise denoising before blending them back together. As a result, all benchmarks are conducted at each dataset’s original aspect ratio and resolution, without additional resizing.
4.2 Comparison
Defocus deblurring.
As stated on Tab. 2, DeblurNet consistently outperforms all baselines across reference and no-reference metrics. On certain NR-IQA measures, it even surpasses the ground-truth. Qualitatively, Fig. 4 (a) shows faithful recovery of the “NEW YORK” text with well-preserved edges. The scene in Fig. 4 (b) is particularly challenging: competing methods recover only a blurry dark stripe and further distort the background, whereas DeblurNet reconstructs structure-consistent content and delivers visually compelling results without introducing artifacts or geometry drift.
Table 4: Quantitative comparison on refocusing benchmark. Results on our LF-Refocus dataset (400 pairs with different focus planes). We compare our DeblurNet+BokehNet pipeline against combinations of existing all-in-focus estimation methods [56, 84] and bokeh synthesis methods [48, 89]. Optimal bokeh level $K$ is determined via per-image binary search following [89]. Our approach achieves superior performance across all metrics, demonstrating the benefit of joint optimization and semi-supervised training.
AiF Estimation Bokeh Synthesis LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ MUSIQ $\uparrow$ NIQE $\downarrow$ CLIP-IQA $\uparrow$
DRBNet [56] BokehMe [48] 0.1471 0.0937 0.9335 43.3087 6.5783 0.4460
BokehDiff [89] 0.1848 0.1030 0.9310 48.0766 5.8195 0.4997
Restormer [84] BokehMe [48] 0.1554 0.0963 0.9303 43.3188 7.1249 0.4782
BokehDiff [89] 0.1860 0.1016 0.9320 48.0326 6.3216 0.5352
DeblurNet (Ours) BokehNet (Ours) 0.1458 0.0850 0.9451 52.6954 5.3343 0.5763
Bokeh synthesis.
Following the BokehDiff protocol [89], we conduct a per-image binary search over the bokeh level $K$ and select the value that maximizes SSIM with the target. All other control signals are held fixed, ensuring a fair comparison in which each method operates at its best attainable setting. As summarized in Tab. 3, our approach attains the highest scores on both fidelity and perceptual metrics, outperforming all baselines. The qualitative results in Fig. 5 show that our method synthesizes bokeh that more closely matches that of real photographs. In the zoomed-in patches, our outputs more faithfully reproduce the ground-truth bokeh than competing baselines. Other methods tend to naively translate the predicted defocus map into blur strength, so any error in the depth scale directly leads to over- or under-blurring. In contrast, BokehNet, trained with our semi-supervised strategy, is more robust to such depth-scale inaccuracies and thus produces a more appropriate amount of blur that better matches the ground truth.
Refocusing.
We build four two-stage baselines by pairing strong AIF estimators (DRBNet [56], Restormer [84]) with bokeh synthesizers (BokehMe [48], BokehDiff [89]), yielding the $2\times 2$ combinations. For each method, we also follow the protocol of BokehDiff [89] to select the best bokeh level $K$ . As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming all baseline pairings. Qualitatively, Fig. 6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content and inaccurate foreground/background separation. In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
@@ 4.2 Comparison
Input GT DRBNet [56] DRBNet [56] Restormer [84] Restormer [84] GenRefocus
+ BokehMe [48] + BokehDiff [89] + BokehMe [48] + BokehDiff [89] (Ours)
@@
Figure 6: Qualitative comparison on refocusing benchmark. Results on LF-Refocus comparing our end-to-end trained pipeline against baseline combinations of existing deblurring [56, 84] and bokeh synthesis [48, 89] methods. Zoomed patches (blue and orange boxes) highlight focal and defocused regions. Baselines struggle with accurate focus-plane placement and produce inconsistent or overly smooth blur. Our method achieves sharp refocusing at the target plane while synthesizing photorealistic bokeh, closely matching ground truth across diverse scenes.
Table 5: Ablation on pipeline design. Comparing one-stage (direct refocusing) and two-stage (deblur then re-bokeh) on LF-Refocus. Both use the same backbone and synthetic data. Two-stage enables better depth control and independent semi-supervised learning per stage, yielding substantial improvements.
@@
One stage 0.1723 0.0976 0.9345
Two stage (ours) 0.1458 0.0850 0.9461
Table 6: Ablation on BokehNet training. Comparing synthetic-only vs. adding real unpaired bokeh images on LF-Bokeh. Second stage uses real bokeh with estimated parameters (Sec. 3.2), initialized from synthetic weights. Real data enables learning authentic optical characteristics beyond simulators.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
Synthetic pair 0.1289 0.0738 0.9461
Synthetic pair+ Real unpaired 0.1047 0.0611 0.9570
4.3 Ablation Study
One stage and two stage refocusing.
We also evaluate a variant that takes an input image and a provided defocus map and directly produces a refocused output in a single pass. Concretely, we instantiate a one-stage model by reusing the BokehNet backbone and training it on the same synthetic pairs, then assess it on the refocusing benchmark. As reported in Tab. 5, the one-stage variant lags the proposed two-stage pipeline by a substantial margin. We attribute this gap to two factors. First, control fidelity: the defocus map relies on depth estimated from a blurry input; without an all-in-focus view, depth quality degrades, leading to misaligned or biased blur fields and weakening refocusing accuracy. Second, data leverage: the one-stage formulation cannot separate the problem into defocus removal and bokeh synthesis, preventing us from injecting real-data supervision tailored to each subtask.
Key idea and BokehNet training ablation.
Our core idea is to use data generation to harvest real-world supervision for training BokehNet, leveraging real unpaired images. To quantify its impact, we evaluate two training settings without introducing additional architectural changes: (i) synthetic pair: BokehNet is trained purely on synthetic paired data; and (ii) synthetic pair + real unpaired: starting from (i), we further train with real unpaired images processed by our pipeline to produce usable supervision. On the bokeh-synthesis benchmark, the synthetic pair setting performs on par with a classical simulator baseline, while the synthetic pair + real unpaired setting delivers substantial improvements across both fidelity and perceptual metrics (Tab. 6).
Application.
As illustrated in Fig. 7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes. Moreover, even though DeblurNet is trained without any text prompts, it can still leverage textual instructions at inference time to refine the all-in-focus output, as shown in Fig. 8. This highlights a key advantage of generative defocus-deblurring models: they naturally support post-hoc, prompt-based correction and editing of the reconstructed content.
@@ 4.3 Ablation Study
Figure 8: Text-guided deblurring. DeblurNet uses optional text prompts to guide reconstruction of severely blurred text. With prompts, we correct hallucinations (red: “DESION” $\rightarrow$ “DESIGN”) and recover accurate text (green: “yeurs” $\rightarrow$ “years”), matching GT. Without guidance, it hallucinates incorrect content.

### v3

@@ 3.3 Bokeh-Shape Aware Synthesis
Table 2: Defocus deblurring benchmark on RealDOF [32] and DPDD [1] datasets. We report reference (LPIPS, DISTS) and no-reference (CLIP-IQA, MANIQA, MUSIQ) perceptual metrics. Best, second best, and third best results are highlighted.
@@
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
Input 0.5241 0.2865 0.3562 0.2213 28.7087 0.3485 0.1827 0.4337 0.3325 45.5376
[…]
INIKNet [47] 0.2851 0.1615 0.3984 0.2580 38.9959 0.1860 0.1235 0.4290 0.3265 47.1961
[…]
Figure 4: Qualitative comparison on defocus deblurring. Visual results on (a) RealDOF [32] and (b) DPDD [1] datasets. Blue boxes on the left indicate cropped regions shown in detail. In (a), most methods fail to resolve the text, whereas both our method and Bokehlicious [57] successfully recover the words “WELCOME” and “OPEN HOURS”. However, Bokehlicious produces noticeable distortions in the right-side characters (e.g., “ME” and “RS”), whereas our approach faithfully restores the original text structure. For the challenging example (b) with severe defocus blur, all baseline methods remain highly blurry and fail to recover meaningful details. In contrast, our method synthesizes a relatively clear and visually compelling output.
3.3.1 Simulator and Data.
Real photographs exhibiting diverse bokeh shapes are rare, and paired AIF–bokeh examples are even scarcer. We thus rely on simulation. A key observation is that when an all-in-focus (AIF) image lacks point-light stimuli, the simulated bokeh carries weak shape cues, making the model reluctant to learn shape-conditioned responses. To address this, we synthesize a PointLight-1K dataset designed to reveal aperture shape. We also extend the classical BokehMe renderer to scatter through a given shape: given an AIF image $I_{\mathrm{aif}}$, depth $D$, focus-plane proxy $D_{\text{focus}}$, bokeh level $K$, and a shape kernel $s$ (binary/raster PSF), the simulator R renders
$\small I_{\mathrm{syn}}\;=\;\mathcal{R}\!\big(I_{\mathrm{aif}},D;\,D_{\text{focus}},K,s\big),$ (6)
@@
3.3.2 Shape-conditioned fine-tuning.
To incorporate the aperture shape $A$, we append its tokens directly to the unified sequence. To ensure that shape edits do not regress the learned bokeh synthesis, we freeze all original LoRA weights and introduce a new, trainable LoRA module. Only this dedicated LoRA is fine-tuned to handle the shape conditioning.
3.4 Pre-deblur Module
We also study a variant that augments DeblurNet with a conservative pre-deblur prior. We train this variant using the same data as the base DeblurNet. Specifically, it condition on $I_{\text{in}}$ and $I_{\text{pd}}$, where $I_{\text{pd}}$ is generated by an off-the-shelf defocus deblurring model. In the implementation, we use a non-generative defocus model [53], whose outputs are typically content-faithful but overly smooth. Qualitative results for this variant are shown in the supplementary material.
3.5 Implementation Details
While our backbone [58] intrinsically supports arbitrary resolutions and aspect ratios without resizing, and our conditioning setup is completely coordinate-independent, ultra-high-resolution inputs (e.g., 4K) still incur prohibitive VRAM consumption and quality degradation. We implement a tiling strategy inspired by [5] during inference. Specifically, we process the image as overlapping patches and blend the tile-wise denoised results. This strategy enables us to process images exactly at their original resolutions, fully preserving image fidelity.
@@ 4.1 Setup
4.1.1 Backbone and training details.
@@
Our backbone is FLUX-1-dev [58], fine-tuned via LoRA [25] with a conditioning scheme following [62, 63]. DeblurNet employs LoRA rank r=128r{=}128, while BokehNet uses r=64r{=}64. Both models are trained with a per-GPU batch size of $1$ and gradient accumulation over $8$ steps on $4{\times}$ RTX A6000 GPUs. DeblurNet is trained for $60\mathrm{K}$ steps. BokehNet is trained in two stages: (i) $40\mathrm{K}$ steps on synthetic data, and (ii) $60\mathrm{K}$ steps on real data.
4.1.2 Runtime and computational cost.
During inference, each stage employs 28 denoising steps. Without the tiling strategy, our pipeline consumes approximately 40 GB of VRAM and takes approximately 50 seconds per image on a single NVIDIA RTX A6000 GPU.
4.1.3 Datasets.
We train DeblurNet on 3.5K pairs sourced from DPDD [1] and a subset of RealBokeh [57]. For BokehNet, to simultaneously maintain realism and controllability, we utilize a hybrid dataset comprising $\sim$70K synthetic pairs derived from [80, 27], combined with approximately 26K real examples sourced from ITW dataset [19], RealBokeh [57], and LFDOF [52].
Table 3: Bokeh synthesis benchmark. Evaluation on LF-Bokeh. Following [87, 43], we use per-image binary search for optimal $K$. Since Bokehlicious [57] uses f-stop rather than $K$ as its control input, we perform a binary search over f-stop values. Due to this fundamentally different control mechanism, we exclude it from the LVCorr evaluation. Note that Bokeh diffusion [19] targets the text-to-image task. To adapt it for image editing, we use textual inversion [61].
Fidelity Controllability
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ LVCorr $\uparrow$
[…]
Bokehlicious [57] 0.1799 0.1062 0.9304 -
[…]
Bokeh Diffusion [19] 0.4999 0.2529 0.7569 0.8954
[…]
@@
Input GT BokehMe Bokehlicious BokehDiff Diffcamera GenRefocus
[43] [57] [87] [69] (Ours)
Figure 5: Qualitative comparison on bokeh synthesis benchmark. Results on LF-Bokeh with zoomed patches (blue and orange boxes) highlighting detail quality. Our method synthesizes bokeh effects that better match ground truth with realistic blur gradients and natural occlusion handling.
4.1.4 Benchmarks and Metrics
We evaluate our method on three tasks using reference-based fidelity and no-reference image quality assessment metrics.
(i) Defocus Deblurring.
We conduct deblurring experiments on RealDOF and DPDD datasets. To assess reconstruction fidelity and perceptual naturalness, we report LPIPS [83], DISTS [16], and no-reference metrics including CLIP-IQA [67], MANIQA [76], and MUSIQ [29].
(ii) Bokeh Synthesis.
We introduce LF-Bokeh, featuring $200$ images with diverse focus planes and aperture sizes synthesized from light-field captures [50, 14]. In real-world scenarios, the absence of ground-truth depth maps and reference bokeh levels $K$ precludes the computation of accurate defocus maps, making the direct evaluation of fidelity and controllability challenging. To assess fidelity, following prior works [87, 43], we conduct a per-image binary search over $K$ and select the value that maximizes SSIM with the target, reporting LPIPS, DISTS, and CLIP-I [48]. To quantify controllability under these constraints, we adopt the LVCorr metric following [80, 19, 69]. Specifically, for images generated from the same all-in-focus input with a fixed focus plane across varying bokeh levels $K$, we compute the Pearson correlation coefficient (CorrCoef) between these $K$ values and the Laplacian variance trend, thereby reflecting the model’s accuracy in rendering the physical blur progression.
(iii) Refocusing.
The LF-Refocus dataset consists of $400$ source–target pairs from LF-Bokeh. Similar to bokeh synthesis, we employ an optimal $K$ search for each pair. We report LPIPS, DISTS, and CLIP-I for fidelity, alongside CLIP-IQA and MUSIQ to evaluate perceptual quality across varying depths of field.
4.2 Comparison
4.2.1 Defocus deblurring.
As reported in Tab. 2, DeblurNet outperforms competing methods across most referenced and non-referenced metrics. While our method yields slightly lower reference scores than Bokehlicious[57] on RealDOF[52], it consistently produces more visually appealing results and demonstrates greater robustness to diverse input conditions, such as grayscale photographs and low-quality, blurry images. Extensive visual comparison supporting this flexibility is provided in the supplementary material. Qualitatively, this perceptual superiority is evident in Fig. 4. As shown in Fig. 4(a), our approach achieves faithful recovery of text details with well-preserved edges. Furthermore, the scene in Fig. 4(b) presents a particularly challenging case: while competing methods recover only a blurry dark stripe and introduce background distortions, DeblurNet reconstructs more consistent content and delivers compelling results.
Table 4: Quantitative comparison on refocusing benchmark. Results on our LF-Refocus dataset. We compare our method against current state-of-the-art refocusing pipelines [69, 64]. Optimal bokeh level $K$ is determined via per-image binary search following [87, 43].Learn2Refocus [64] takes a specific focal position as the input condition to synthesize the remaining images in the focal stack. To ensure coverage of the full range of focal planes, we perform inference twice separately: once conditioned on the minimum focal position (1) and once on the maximum (9). We then evaluate all outputs from both generated stacks and report the one yielding the highest SSIM.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$ CLIP-IQA $\uparrow$ MUSIQ $\uparrow$
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input GT DiffCamera Learn2Refocus GenRefocus
[69] [64] (Ours)
@@ 4.2 Comparison
Figure 6: Qualitative comparison on the refocusing benchmark. Results on the LF-Refocus dataset with zoomed patches (blue and orange boxes) highlight the fidelity of focal and defocused regions. We observe that DiffCamera [69] struggles to reconstruct sharp structural details due to its strict input conditions and often assigns incorrect blurriness levels to background regions. Similarly, Learn2Refocus [64] fails to achieve accurate focus-plane placement, resulting in unintended blur of the target object. In contrast, our method achieves precise refocusing at the target plane while synthesizing photorealistic bokeh, closely matching the ground truth across diverse scenes.
4.2.2 Bokeh synthesis.
As presented in Tab. 3, our approach outperforms all other methods across fidelity metrics. In terms of controllability, which measures how well the generated bokeh aligns with the conditioned bokeh level $K$, BokehMe [43] achieves near-perfect control owing to its hybrid design that incorporates a classical, physically motivated renderer. Nevertheless, our method outperforms all diffusion-based baselines, demonstrating that our learning strategy effectively preserves both realism and accuracy. Furthermore, it is worth noting the specific limitations of competing methods: DiffCamera [69] requires inputs with a fixed resolution and aspect ratio due to its reliance on explicit focus point conditioning, rendering it incompatible with tiling strategies; consequently, resizing operations severely degrade its controllability. Meanwhile, Bokeh Diffusion [19] relies on an inversion [61] process that not only prolongs inference time but also strips the image of its original high-frequency details, thereby lowering overall fidelity. The qualitative results shown in Fig. 5.
4.2.3 Refocusing.
As shown in Tab.4, our approach achieves the best results on both fidelity and perceptual metrics, outperforming other baselines. Qualitatively, Fig. 6 shows that competing methods tend to produce spatially ambiguous blur and fail to place the focal plane correctly, resulting in uniformly soft content. In contrast, our approach restores fine details at the intended focus while synthesizing physically consistent bokeh elsewhere.
4.3 Ablation Study
Table 5: Comparison of pipeline designs. Comparing direct refocusing with our proposed deblurring, then the bokeh synthesis approach on LF-Refocus.
@@ 4.3 Ablation Study
Direct refocusing 0.1723 0.0976 0.9345
Deblurring then bokeh synthesis (Ours) 0.1324 0.0759 0.9571
Table 6: Ablation on BokehNet training datasets. Comparing combinations of datasets: (a) Synthetic data, (b) ITW dataset[19], and (c) LFDOF [52] and RealBokeh[57] on LF-Bokeh. The results show that combining all three yields the best optical characteristics.
Datasets Metrics
(a) Synthetic (b) ITW (c) LFDOF & RealBokeh LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-I $\uparrow$
- - 0.1289 0.0738 0.9461
- 0.1156 0.0665 0.9529
[…]
4.3.1 Direct refocusing vs. deblurring then bokeh synthesis.
We evaluate a baseline that directly predicts a refocused image from a blurry input and a defocus map. Trained on identical synthetic data using the BokehNet backbone, this direct approach significantly underperforms our two-stage pipeline (Tab. 5). We attribute this gap to two factors: (1) control fidelity: estimating depth from blurry inputs degrades defocus map accuracy, causing misaligned blur fields; and (2) data leverage: a direct formulation precludes injecting tailored real-data supervision into separate deblurring and synthesis subtasks.
4.3.2 BokehNet training ablation.
To quantify the impact of different training data distributions, we evaluate BokehNet training under various dataset combinations on LF-Bokeh. We consider three primary data sources: (a) synthetic data, (b) ITW dataset[19], and (c) LFDOF [52] and RealBokeh [57]. As shown in Tab. 6, a baseline trained purely on (a) provides a foundational performance but leaves room for improvement. Introducing either (b) or (c) to the synthetic baseline noticeably enhances performance. Ultimately, the comprehensive setting that combines all three sources yields the most substantial improvements, delivering the best optical characteristics and real-world generalization.
4.3.3 Aperture shape control.
As illustrated in Fig. 7, after we fine-tune BokehNet on our point-light 1K dataset, the model can synthesize bokeh with diverse, user-specified aperture shapes.
@@
Figure 8: Failure case: transparent surfaces. With a transparent glass, focusing on the foreground may fail to correctly blur the background due to unreliable defocus maps.
Figure 9: Failure case: hallucinated details. With severely blurred inputs, our model may hallucinate incorrect results.